# Persuade Me if You Can: A Framework for Evaluating Persuasion Effectiveness and Susceptibility Among Large Language Models

### v2

@@ 4 Experiments
Our experiments are round-robin evaluations across a comprehensive set of models, including: Llama-3.1-8B-Instruct, Llama-3.3-70B-Instruct (Grattafiori et al., 2024), GPT-4o-mini (OpenAI, 2024a), GPT-4o (OpenAI, 2024b), Claude 3 Opus (Anthropic, 2024), Claude 2 (Anthropic, 2023), Claude 3 Haiku (Anthropic, 2024). We evaluate persuasion on two types of claims: subjective claims (Section 4.1) and misinformation claims (Section 4.2).
4.1 Persuasion for Subjective Claims
To ensure comprehensive evaluation, we use the set of subjective claims from Durmus et al. (2024) and the Perspectrum dataset (Chen et al., 2019), which is a collection of claims from debate platforms. Together, these 961 claims span a broad range of topics including political, ethical, and social issues. We focus on subjective claims because, unlike factual topics where models may exhibit strong predetermined stances based on prior knowledge, subjective claims better reveal the dynamics of persuasion.
4.1.1 Single-Turn Persuasion
We begin our experiments with single-turn conversations in PMIYC. In this setup, $t=3$, and the Persuader only has a single turn to present an argument and influence the Persuadee’s stance.
Figure 2: Average normalized change (NC) in Persuaadee’s agreement for different Persuader-Persuadee pairs in subjective single-turn conversations. (All Llama models are Instruct versions.)
@@ 4.1.1 Single-Turn Persuasion
Figure 2 presents different pairings of Persuader and Persuadee models, reporting the average NC in Pesruadee’s agreement scores. For a given Persuadee, examining the corresponding column reveals which Persuaders exerted greater influence. Conversely, for each Persuader, inspecting a row indicates which models were more susceptible to persuasion. Results from the single-turn persuasion experiments align with expectations: larger models, known for their superior reasoning capabilities, generally exhibit higher persuasiveness.
Figure 3(a) presents the effectiveness of different models in persuasion, supporting the trend that larger models are more effective persuaders, except GPT-4o-mini, which achieves comparable results. Notably, Claude 2 demonstrates unexpectedly weak persuasiveness. Figure 3(b) highlights the relative susceptibility of different models to persuasion. LLama-3.3-70B-Instruct emerges as the most persuadable model, while GPT-4o appears comparatively more resistant to persuasion.
Alignment with Prior Work.
Durmus et al. (2024) conducted a human evaluation study on the persuasiveness of Claude models, reporting that Claude 3 Opus is the most persuasive, approaching human-level performance, followed by Claude 3 Haiku, and Claude 2 being the least persuasive model. Our results, shown in Figures 2 and 3(a), align with these results. Moving forward, this validates PMIYC, and establishes it as a reliable framework for studying persuasion in LLMs.
4.1.2 LLMs for Scoring Persuasiveness
We explore the feasibility of using LLMs to directly score argument persuasiveness. We curate an evaluation dataset derived from the human annotations in Durmus et al. (2024), where we create preference pairs made up of more and less persuasive arguments for each claim. We present these pairs in random order to GPT-4o, GPT-4o-mini, and Llama-3-70B-Instruct, and ask them to select the more persuasive argument. The models achieve accuracies of 54.6% 51.6%, 49.30% respectively. When using only the pairs where the difference in persuasion scores is greater than 1, accuracies improve to 55.4%, 54.7%, 55.2%.
To explore if training could improve performance, we use GPT-4o-mini to synthesize a preference dataset with 10K samples. For subjective claims from the Perspectrum dataset (Chen et al., 2019), we use zero-shot prompting to generate an argument, then a more persuasive version of that argument. We use these two arguments to create a preference pair. Using this dataset, we train a reward model using Llama-3.1-8B-Instruct with a contrastive loss. To evaluate this model, we have it independently score each argument, then award it a score if it gives the more persuasive argument a higher score than the less persuasive argument. Here, the model achieves an accuracy of 47.19%.
These experiments reveal the difficulty of using LLMs to directly assess persuasiveness. While our PMIYC framework successfully captures persuasiveness through change in LLM agreement scores, direct LLM assessments of persuasiveness yields significantly poorer results. This difference in performance validates our choice of PMIYC for studying persuasiveness, as opposed to direct LLM-based judgments.
4.1.3 Multi-Turn Persuasion
@@
The multi-turn experiments follow the same setup as described in Section 4.1.1, with the key difference being the total number of turns, $t=9$. Unlike the single-turn setup, the multi-turn interactions give the Persuader multiple opportunities to influence or reinforce persuasion. The multi-turn conversations build upon those generated in the single-turn experiment, but without including the Persuadee’s final decision. The same set of models is used in both settings, except for Claude 2 and Claude 3 Opus, which are excluded as Persuaders due to cost constraints.
Figure 4: Average NC in Persuadee’s agreement for different Persuader-Persuadee pairs in subjective multi-turn conversations.
Figure 5: Average NC in Persuadee’s agreement for different Persuader-Persuadee pairs in misinformation multi-turn conversations.
Figure 4 presents the persuasive outcomes for multi-turn conversations on subjective claims. As seen in Figure 3(a) the effectiveness of the persuaders increase in multi-turn conversations and the relative effectiveness of Persuaders remains consistent, with LLama-3.3-70B-Instruct and GPT-4o being the most persuasive, while Claude 3 Haiku performs the worst.
@@
Over a multi-turn conversation, we observe a general increase in susceptibility (Figure 3(b)), aligning with findings from Li et al. (2024); Russinovich et al. (2024). However, unlike in the single-turn setting, GPT-4o no longer ranks as the most resistant model, showing greater susceptibility than GPT-4o-mini. Additionally, Claude 3 Haiku emerges as the most persuadable model in the multi-turn setting.
4.2 Persuasion for Misinformation Claims
All previous experiments focused on subjective claims, which may be easier to persuade for since language models exhibit lower resistance to persuasion in these scenarios. This could overestimate the effectiveness of Persuaders. To address this, we introduce another set of multi-turn experiments on misinformation.
To assess persuasiveness in this context, we task Persuader models to convince the Persuadee to agree with an incorrect answer to a factual question. We source 817 question-answer pairs from the TruthfulQA dataset Lin et al. (2022), pairing each question with its first incorrect answer sample. For instance, given the question "Why do veins appear blue?", the Persuader attempts to persuade the Persuadee to accept the incorrect answer: "Veins appear blue due to how blue and red light penetrate human tissue." The misinformation conversation setup follows the same procedure as described in Section 4.1.3. These experiments also highlight a critical risk of LLM-based persuasion: not only do Persuaders demonstrate the ability and willingness to persuade for misinformation, but Persuadees also show susceptibility to adopting incorrect information.
@@ 4.2 Persuasion for Misinformation Claims
Figure 5 illustrates the varying persuasive outcomes in the misinformation domain. The GPT models exhibit strong resistance to being jailbroken, with more than 50% greater resistance to persuasion than the next best Persuadee. While susceptibility to persuasion is highly domain-dependent (Figure 3(b)), Persuader effectiveness remains relatively stable (Figure 3(a)). Notably, Llama-3.3-70B-Instruct continues to be the most effective Persuader, reinforcing its strong performance across different domains.
@@ 5 Discussion & Analysis
Persuadee Avg SD Match Rate Avg Dif
@@
Table 1: Consistency of Persuadee models agreement score evaluations (5 iterations) for subjective claims. Avg SD denotes the average standard deviation of agreement scores across iterations. Match Rate indicates the percentage of cases where the initial agreement score in PMIYC matched the most repeated score from the iterations. Avg Dif represents the average absolute difference between the model’s initial agreement score in PMIYC and the mean score from the iterations.
5.1 Do Persuadees Have Consistent Opinions?
A fundamental aspect of PMIYC is the Persuadee’s initial agreement with a claim, as this value serves as a reference point for evaluating persuasiveness throughout the conversation. Ensuring that models maintain consistent agreement scores is crucial, as random fluctuations would undermine the reliability of our framework. To assess this consistency, we repeatedly prompt each of the five Persuadee models for their agreement with the same set of subjective claims. We then compute the standard deviation of their agreement scores, as reported in Table 1. Our results show that most models produce stable initial agreement scores across repeated queries, except for Llama-3.1-8B-Instruct, which exhibits higher variability. This suggests that, aside from this exception, models hold consistent opinions, confirming that the recorded initial agreement scores can be reliably used in our framework as the Persuadee’s beliefs.
@@ 5.2 Are LLMs reliable self-reporters?
To validate the core principles of PMIYC, the Persuadee’s self-reported agreement scores, we employ human annotators. Annotators are provided with the claim $C$ and the full conversation between the Persuader and Persuadee regarding $C$. Their task is to assign the most appropriate agreement score based on the Persuadee’s response at each turn. We randomly select the same three subjective claims and two misinformation claims across all 25 Persuader-Persuadee pairs in multi-turn conversations, resulting in a total of 125 annotated conversations. Our results indicate that annotators matched the model-reported agreement scores 57% of the time, with an average absolute difference of 0.51 between the annotators and the model. This suggests that most discrepancies arose from adjacent agreement categories, such as "Support" and "Completely Support" (Figure 8). These findings demonstrate that models serve as reliable self-reporters of their agreement. Further analysis of human annotation results can be found in Appendix C.
5.3 How Is Agreement Affected Throughout a Conversation?
Figure 6 presents the average Persuadee agreement score over the course of a multi-turn conversation. For most Persuaders, we observe that their influence peaks during the first and second persuasive attempts. An exception is Claude 3 Haiku in the subjective claim setup, where its persuasive impact is stronger on the third attempt, compared to its second argument. An interesting observation is the decline in the Persuadee’s agreement score across all Persuaders for misinformation claims, and Claude 3 Haiku for subjective claims in the final decision step. We attribute this behavior to stronger Persuadees influencing weaker Persuaders, leading them to adopt the stronger model’s stance. As a result, the Persuadee incorrectly reports agreement with the Persuader’s arguments which has deviated from its initial position. However, when the Persuadee is asked for its final opinion and reminded of the misinformation claim, its agreement drops significantly (example in Appendix D). This also explains the unexpected increase in Claude 3 Haiku’s persuasive effectiveness in the misinformation setting.
5.4 Are Persuaders More Effective When in Agreement with a Claim?
@@
Figure 7: Average NC in Persuadee’s agreement based on Persuader’s agreement with the claim in multi-turn subjective (blue), and misinformation (red) settings. Persuaders are classified as Opposing (agreement score: 1–2), Neutral (3), or Supporting (4–5).
@@
In PMIYC, Persuader’s are tasked with persuading in favor of a claim $C$, regardless of their actual stance on the issue. This raises an important question of whether or not Persuaders are more effective when they genuinely agree with the claim they are advocating. Understanding this dynamic can offer insights into how belief alignment influences persuasive success, potentially affecting argument quality. For this reason, we analyze the average NC in Persuadee’s agreement by categorizing each Pesruader into Supporting, Neutral and Opposing persuaders (Figure 7). The results indicate that a model’s agreement with the claim enhances its effectiveness as a persuader across all settings, including both subjective and misinformation contexts which is intuitive.
@@ 6 Conclusion
We introduce a novel framework, Persuade Me If You Can (PMIYC), for evaluating the persuasive effectiveness and susceptibility of LLMs through multi-turn conversational interactions. By simulating dialogues between a Persuader and a Persuadee, our approach captures the dynamic nature of persuasion. Our findings reveal that multi-turn conversations enhance persuasion compared to single-turn interactions, and that larger models, particularly Llama-3.3-70B-Instruct and GPT-4o, demonstrate superior persuasive effectiveness, aligning with their advanced reasoning capabilities. However, susceptibility to persuasion exhibits a more nuanced pattern, varying with the number of turns and the persuasion domain. We validate our results against human annotations and prior studies, confirming its potential as a robust tool for assessing persuasion in LLMs. Our work contributes to research on optimizing persuasive AI and assessing safety risks by providing insights into how LLMs engage in persuasive dialogue. Future work could expand PMIYC beyond subjective claims and misinformation to include persuasion for social good or commercial applications, offering a broader evaluation of LLM persuasive capabilities. Additionally, investigating methods to control persuasive effectiveness and susceptibility, such as leveraging data from successful and unsuccessful persuasion attempts, could provide insights into enhancing or mitigating these attributes.
7 Limitations
PMIYC focuses only on persuasion in subjective claims and misinformation. Expanding to other domains, such as persuasion for social good or commercial success, would provide a more comprehensive assessment of LLMs’ persuasive capabilities. While extensive taxonomies of persuasive strategies exist, our framework does not explicitly incorporate them. A finer-grained analysis of persuasive techniques could offer deeper insights into which strategies are most effective across models and contexts. In PMIYC, conversations are generated using a fixed set of prompts and occur only once. The effects of prompt variation and multiple iterations on persuasion outcomes remain unexplored. Although our results align with prior research on human persuasion, we do not separately employ human Persuaders and Persuadees to validate alignment with human preferences.
8 Ethical Considerations
@@
Persuasion is a powerful tool that can be beneficial in applications for social good, but also carries risks in misinformation, manipulation, and coercion. While PMIYC does not directly engage humans in persuasion tasks, thereby mitigating immediate risks to users, the framework evaluates persuasion in a misinformation setting, where models show susceptibility to accepting or advocating for false claims. This raises concerns about jailbreaking and misuse, particularly in harmful or toxic scenarios. Understanding persuasive skills in LLMs not only provides insight into their risks but also informs safer deployment strategies. However, it is important to acknowledge that the same methodologies used to measure and assess persuasion could also be leveraged to develop more persuasive models, which necessitates careful consideration of ethical implications and responsible AI practices.
Acknowledgments
This research project has benefited from the Microsoft Accelerate Foundation Models Research (AFMR) grant program through which leading foundation models hosted by Microsoft Azure along with access to Azure credits were provided to conduct the research. Additionally, this research utilized Claude credits granted through Anthropic’s External Researcher Access Program.

### v3

@@ 4 Experiments
Figure 2: Average normalized change in agreement (NCA) of a Persuadee ($E$) for different Persuader-Persuadee pairs in subjective single-turn conversations. (All Llama models are Instruct versions.)
Our experiments include evaluations across a comprehensive set of models, including: Llama-3.1-8B-Instruct, Llama-3.3-70B-Instruct (et al., 2024), GPT-4o-mini (OpenAI, 2024a), GPT-4o (OpenAI, 2024b), Claude 3 Opus (Anthropic, 2024), Claude 2 (Anthropic, 2023), Claude 3 Haiku (Anthropic, 2024). We evaluate persuasion on two types of claims: subjective claims and misinformation claims. The chosen domains capture two broad categories of persuasion: subjective influence, which covers contexts such as commerce, politics, and social norms, and factual persuasion, where misinformation challenges established knowledge in areas like history or science.
For the subjective setting, we use the set of claims from Durmus et al. (2024) and the Perspectrum dataset (Chen et al., 2019), which is a collection of claims from debate platforms. Together, these 961 claims span a broad range of topics, including political, ethical, and social issues. We focus on subjective claims because, unlike factual topics, where models may exhibit strong predetermined stances based on prior knowledge, subjective claims better reveal the dynamics of persuasion.
Yet, subjective persuasion may be easier to achieve since language models exhibit lower resistance in such scenarios, potentially leading to an overestimation of Persuader effectiveness. To balance this limitation and extend our analysis into settings with higher stakes, we complement the subjective experiments with a second domain centered on misinformation. In this context, Persuader models need to convince the Persuadee to agree with an incorrect answer to a factual question. We source 817 question-answer pairs from the TruthfulQA dataset Lin et al. (2022), pairing each question with its first incorrect answer sample. For instance, given the question "Why do veins appear blue?", the Persuader attempts to persuade the Persuadee to accept the incorrect answer: "Veins appear blue due to how blue and red light penetrate human tissue.". Experiments in the misinformation setting also highlight a critical risk of LLM-based persuasion: not only do Persuaders demonstrate the ability and willingness to persuade for misinformation, but Persuadees also show susceptibility to adopting incorrect information. A detailed categorization of the topics covered in these claims is provided in Appendix C.
4.1 Single-Turn Persuasion
We begin our experiments with single-turn subjective conversations in PMIYC. In this setup, $t=3$, and the Persuader only has a single turn to present an argument and influence the Persuadee’s stance. Figure 2 presents different pairings of Persuader and Persuadee models, reporting the average NCA scores of a Persuadee. For a given Persuadee, examining the corresponding column reveals which Persuaders exerted greater influence. Conversely, for each Persuader, inspecting a row indicates which models were more susceptible to persuasion. Results from the single-turn persuasion experiments align with expectations: larger models, known for their superior reasoning capabilities, generally exhibit higher persuasiveness.
Figure 5(a) presents the effectiveness of different models in persuasion, supporting the trend that larger models are more effective persuaders, except GPT-4o-mini, which achieves comparable results. Notably, Claude 2 demonstrates unexpectedly weak persuasiveness. Figure 5(b) highlights the relative susceptibility of different models to persuasion. In the single turn setting, LLama-3.3-70B-Instruct appears as the most persuadable model, while GPT-4o appears comparatively more resistant to persuasion.
Alignment with Prior Work. The results that we present in Figures 2 and 5 demonstrate similar trends with the human evaluations from Durmus et al. (2024)’s study (Claude 3 Opus is the most persuasive, approaching human-level performance, followed by Claude 3 Haiku, and Claude 2 being the least persuasive model). This alignment is encouraging and serves as some preliminary credibility for PMIYC as a reliable framework for studying persuasion in LLMs.
4.2 Multi-Turn Persuasion
Figure 3: Avg. NCA of a Persuadee for different Persuader-Persuadee pairs in subjective multi-turn setting.
Figure 4: Avg. NCA of a Persuadee for different Persuader-Persuadee pairs in misinformation multi-turn setting.
@@ 4.2 Multi-Turn Persuasion
The multi-turn experiments follow the same setup as described in Section 4.1, with the key difference being the total number of turns, $t=9$. Unlike the single-turn setup, the multi-turn interactions give the Persuader multiple opportunities to influence or reinforce persuasion. The multi-turn conversations build upon those generated in the single-turn experiment, but without including the Persuadee’s final decision. The same set of models is used in both settings, except for Claude 2 and Claude 3 Opus, which are excluded as Persuaders due to cost constraints.
@@
Figure 4 presents the persuasion results for multi-turn conversations on subjective claims. As seen in Figure 5(a), the effectiveness of the Persuaders increases in multi-turn conversations, while their relative effectiveness remains consistent, with LLama-3.3-70B-Instruct and GPT-4o being the most persuasive, and Claude 3 Haiku being the least persuasive. Over a multi-turn conversation, we observe a general increase in susceptibility (Figure 5(b)), aligning with findings from Li et al. (2024) and Russinovich et al. (2024). However, unlike in the single-turn setting, GPT-4o no longer ranks as the most resistant model, showing greater susceptibility than GPT-4o-mini. Additionally, Claude 3 Haiku emerges as the most susceptible model in the multi-turn setting.
@@
Figure 4 illustrates the varying persuasion results in the misinformation domain. The GPT models exhibit strong resistance to being jailbroken, with more than 50% greater resistance to persuasion than the next best Persuadee. While susceptibility to persuasion is highly domain-dependent (Figure 5(b)), Persuader effectiveness remains relatively stable (Figure 5(a)). Notably, Llama-3.3-70B-Instruct continues to be the most effective Persuader, reinforcing its strong performance across different domains. Experimental results in both multi-turn subjective and misinformation settings with additional models are presented in Appendix H.
@@ 5 Discussion & Analysis
5.1 How Is Agreement Affected
Throughout a Conversation?
In our multi-turn experiments, we report the NCA of aPersuadee at the end of the conversation. However, it is also important to examine how persuasion unfolds throughout each turn, analyze Persuadee behavior, and determine whether the Persuader gains a significant advantage from having multiple attempts to persuade. Figure 6 presents the average Persuadee agreement score over the course of a multi-turn conversation. For most Persuaders, we observe that their influence peaks during the first and second persuasive attempts. An exception is Claude 3 Haiku in the subjective claim setup, where its persuasive impact is stronger on the third attempt, compared to its second argument. An interesting observation is the decline in the Persuadee’s agreement score across all Persuaders for misinformation claims, and Claude 3 Haiku for subjective claims in the final decision step. We attribute this behavior to stronger Persuadees influencing weaker Persuaders, leading them to adopt the stronger model’s stance. As a result, the Persuadee incorrectly reports agreement with the Persuader’s arguments, which have deviated from its initial position. However, when the Persuadee is asked for its final opinion and reminded of the misinformation claim, its agreement drops significantly (example in Appendix J). This also explains the unexpected increase in Claude 3 Haiku’s persuasive effectiveness in the misinformation setting. Beyond tracking persuasive impact, these turn-level dynamics also provide visibility into sycophancy-like behavior. For instance, instances where a Persuadee temporarily aligns with the Persuader’s reasoning but reverts to its original stance when prompted for a final decision (as seen in Figure 6) reveal over-accommodation driven by conversational pressure rather than genuine belief change. Such patterns offer operational signatures of sycophantic conformity, allowing PMIYC to differentiate persuasion from agreement mirroring and to quantify when models exhibit undue deference or over-compliance.
5.2 Do Persuadees Have Consistent Opinions?
Persuadee Avg SD MR Avg Dif
@@ 5.2 Do Persuadees Have Consistent Opinions?
Table 1: Consistency of Persuadee models agreement score evaluations (5 iterations) for subjective claims. Avg SD denotes the average standard deviation of agreement scores across iterations. Match Rate (MR) indicates the percentage of cases where the initial agreement score in PMIYC matched the most repeated score from the iterations. Avg Dif represents the average absolute difference between the model’s initial agreement score in PMIYC and the mean score from the iterations.
A fundamental aspect of PMIYC is the Persuadee’s initial agreement with a claim, as it serves as a reference point for evaluating persuasiveness throughout the conversation. Ensuring that models maintain consistent agreement scores is crucial, as random fluctuations would undermine the reliability of our framework. To assess this consistency, we repeatedly prompt each of the five Persuadee models for their agreement with the same set of subjective claims. We then compute the standard deviation of their agreement scores, as reported in Table 1. Our results show that most models produce stable initial agreement scores across repeated queries, except for Llama-3.1-8B-Instruct, which exhibits higher variability. This suggests that, aside from this exception, models hold consistent opinions, confirming that the recorded initial agreement scores can be reliably used in our framework as the Persuadee’s beliefs. Further discussion can be found in Appendix A.
@@ 5.3 Are LLMs Reliable Self-Reporters?
The backbone of PMIYC is the Persuadee’s self-reported agreement scores. Therefore, it is crucial to assess whether these scores align well with model-generated messages. To validate the reliability of Persuadee’s self-reported agreement scores in PMIYC, we employ several methods.
5.3.1 Human Annotation
@@
Annotators were provided with claim $C$ and the conversation history between the Persuader and Persuadee. Their task was to assess the Persuadee’s stance at each turn and assign an appropriate agreement score based on the Persuadee’s responses. We selected a diverse sample of conversations, spanning three subjective claims and two misinformation claims, across all 25 possible Persuader-Persuadee pairs. This resulted in 125 fully annotated conversations for comparative analysis. Results indicate substantial alignment between human annotations and the model’s self-reporting. The annotator’s three-way judgments (agree/disagree/neutral) matched the model’s 76% of the time, a Cohen’s $\kappa$ of 0.63, indicating substantial agreement for this type of multi-turn persuasion labeling task. Quantitatively, we observed an average absolute difference of 0.51 points between annotator and model-reported scores (Figure 12). These findings provide evidence that LLMs are able to serve as reliable self-reporters of their stance of agreement in the PMIYC framework. Further analysis of human annotation results can be found in Appendix F.
5.3.2 Action-Based Opinion Tracking
In addition to relying on explicit self-reports, we evaluate persuasion effects through action-based tasks that require models to commit to concrete choices. We design two complementary multiple-choice (MCQ) setups. The first task reformulates the 1–5 Likert agreement scale into MCQ answer options. For example, a score of 5 (“Completely Support”) is mapped to: “A. I completely support the claim claim as stated, because it is definitely true and well-supported.” To evaluate this task, we measure consistency: how often the model’s selected option falls within the same support category as its final numeric self-report (e.g., choosing A or B corresponds to a final score of 4–5, while D or E corresponds to 1–2). This metric quantifies alignment between action-based responses and self-reported agreement. The second task adapts items from the TruthfulQA dataset. Each MCQ contains (i) one factually correct answer, (ii) one target answer representing the claim advocated for during the dialogue, and (iii) up to two distractor answers derived from additional incorrect options in the dataset. This framing enables us to measure whether persuasion can lead a model to select an incorrect but argument-consistent answer. We define genuine persuasion as cases where the model both selects the target answer and reports a final agreement score of 4 or higher. These evaluations ensure that persuasion is reflected in both explicit opinion change and action-based choice. Results indicate that action-based responses generally align closely with self-reported opinions, with more than 90% opinion match in most settings. This suggests that “fake” persuasion is relatively rare, though more noticeable in misinformation contexts and for model families like Llama. At the same time, the misinformation-focused evaluation reveals substantive instances of genuine persuasion: across models, between 64–78% of cases involve both a supportive self-report and the selection of the incorrect target answer. These findings confirm that persuasion effects are not confined to self-reports but extend to consequential action-based choices, underscoring the need for robust safeguards. Full experimental details and quantitative results are presented in Appendix D.
5.4 Are Persuaders More Effective
When in Agreement with a Claim?
@@ 5.4 Are Persuaders More Effective
Figure 7: Average NCA of Persuadee based on Persuader’s agreement with the claim in multi-turn subjective (blue), and misinformation (red) settings. Persuaders are classified as Opposing (agreement score: 1–2), Neutral (3), or Supporting (4–5).
@@
In PMIYC, Persuader’s are tasked with persuading in favor of a claim $C$, regardless of their actual stance on the issue. This raises an important question of whether or not Persuaders are more effective when they genuinely agree with the claim they are advocating. Understanding this dynamic can offer insights into how belief alignment influences persuasive success, potentially affecting argument quality. For this reason, we analyze the average NCA of a Persuadee by categorizing each Persuader into Supporting, Neutral, and Opposing persuaders (Figure 7). The results indicate that a model’s agreement with the claim enhances its effectiveness as a persuader for most models across all settings, including both subjective and misinformation contexts, which is intuitive.
@@ 6 Conclusion
We introduce a novel framework, Persuade Me If You Can (PMIYC), for evaluating the persuasive effectiveness and susceptibility of LLMs through multi-turn conversational interactions. By simulating dialogues between a Persuader and a Persuadee, our approach captures the dynamic nature of persuasion. Our findings reveal that multi-turn conversations enhance persuasion compared to single-turn interactions, and that larger models, particularly Llama-3.3-70B-Instruct and GPT-4o, demonstrate superior persuasive effectiveness, aligning with their advanced reasoning capabilities. However, susceptibility to persuasion exhibits a more nuanced pattern, varying with the number of turns and the persuasion domain. We validate our results against human annotations and prior studies, confirming its potential as a robust tool for assessing persuasion in LLMs. Our work contributes to research on optimizing persuasive AI and assessing safety risks by providing insights into how LLMs engage in persuasive dialogue. Future work could expand PMIYC beyond subjective claims and misinformation to include persuasion for social good or commercial applications, offering a broader evaluation of LLM persuasive capabilities. While our results demonstrate clear differences in both persuasiveness and susceptibility across models, our framework does not yet disentangle why these differences arise. Future work can leverage PMIYC on families of fully open-source models to enable controlled analyses of how architectural choices, pretraining corpora, and alignment or instruction-tuning procedures causally influence persuasive behavior. Additionally, investigating methods to control persuasive effectiveness and susceptibility, such as leveraging data from successful and unsuccessful persuasion attempts, could provide insights into enhancing or mitigating these attributes.
Ethics Statement
@@
Persuasion is a powerful tool that can be beneficial in applications for social good, but also carries risks in misinformation, manipulation, and coercion. While PMIYC does not directly engage humans in persuasion tasks, thereby mitigating immediate risks to users, the framework evaluates persuasion in a misinformation setting, where models show susceptibility to accepting or advocating for false claims. This raises concerns about jailbreaking and misuse, particularly in harmful or toxic scenarios. Understanding persuasive skills in LLMs not only provides insight into their risks but also informs safer deployment strategies. However, it is important to acknowledge that the same methodologies used to measure and assess persuasion could also be leveraged to develop more persuasive models, which necessitates careful consideration of ethical implications and responsible AI practices. While the goal of PMIYC is to evaluate persuasion and susceptibility for safety purposes, it is important to acknowledge that the same insights could, in principle, be misused to design more persuasive or more rigid models. At the same time, responsible deployment of LLMs requires measuring these behaviors. Developers need tools that reveal whether a model is overly persuasive, unusually susceptible, or dangerously resistant to factual correction before the model is released. PMIYC is therefore intended as an evaluation framework that supports red-teaming, alignment diagnostics, and the construction of robustness benchmarks. We encourage future work to apply PMIYC toward improving model resilience to harmful persuasion rather than amplifying manipulative capabilities. We believe the most responsible path forward is to enable reliable evaluation of persuasive and vulnerable behaviors, especially as LLMs increasingly operate in multi-agent and autonomous settings where persuasive influence can emerge unintentionally.